A crane thatlifts a60kg mss 100m in60s calculate power

1
Expert's answer
2022-04-12T11:30:25-0400
"P={\rm\frac{work\; done}{time}}=\frac{mgh}{t}"

"P=\frac{60*9.8*100}{60}=980\:\rm W"